OpenStudy (anonymous):

Help ASAP!! these are homework questions please help!!! I will fan and medal!!

OpenStudy (morganchavez):

ok what do u need

OpenStudy (anonymous):

Compare the two numbers, using >, <, or =: -8 and -3. (Points : 1) -8 > -3 -8 < -3 -8 = -3 Question 2. 2. Place the numbers in order from least to greatest: 0, 8, -5, 6, -9. (Points : 1) -9, -5, 0, 6, 8 -5, -9, 0, 6, 8 0, -5, 6, 8, -9 Question 3. 3. Five people are playing a game where the scores can be positive or negative. The winner is the person with the lowest score. These are the current scores: Mandy 0, Mark 5, Ted -2, Sandy -10, Otis 8. Who is winning? (Points : 1) Mandy Mark Ted Sandy Otis Question 4. 4. Match each of the subtraction problems with the problem changed to addition. (Points : 1) Potential Matches: 1 : 5 + -14 2 : -5 + -14 3 : 5 + 14 4 : -5 + 14 Answer : -5 - 14 = : 5 - 14 = : -5 - -14 = : 5 - -14 = Question 5. 5. 500 + -500 = (Points : 1) 1,000 0 -1,000 Question 6. 6. 40 + -23 = (Points : 1) -63 63 17 -17 Question 7. 7. -15 + -13 = (Points : 1) -28 28 -2 2 Question 8. 8. In a game where you can earn positive and negative points, you made these points on the first 5 plays: +10, -42, +39, -25, -5. What is your current number of points in the game? (Points : 1) -23 23 121 -121 Question 9. 9. 9 + (-10) + (-20) = (Points : 1) 39 -39 21 -21 Question 10. 10. 5 - 20 = (Points : 1) 25 -25 -15 15 Question 11. 11. 18 - 12 = (Points : 1) 6 -6 30 -30 Question 12. 12. -6 - 7 = (Points : 1) -1 1 -13 13

OpenStudy (anonymous):

ok

OpenStudy (morganchavez):

holyy jebs ok how long u got ? time wise

OpenStudy (morganchavez):

cuz u gnna have to give me a minute

OpenStudy (anonymous):

that's totaly fine.. take all the time u need!!

OpenStudy (morganchavez):

ok lol

OpenStudy (anonymous):

@morganchavez ??????

OpenStudy (morganchavez):

hang on

OpenStudy (anonymous):

okay:) @morganchavez

OpenStudy (morganchavez):

#1:B #2:A #3:E #4: 1:B,2:A,3:D,4:C #5:B #6:C #7:A #8:A #9:D #10:C #11:A #12:C

OpenStudy (morganchavez):

there you go @lee_actress

OpenStudy (anonymous):

Thank you so much!! Thank you so so so much!!1

OpenStudy (morganchavez):

no problem :)